The Devon Motor Company produces automobiles. On April 1st the company had no beginning inventories and it purchased 6,950 batteries at a cost of $125 per battery. It withdrew 6,400 batteries from the storeroom during the month. Of these, 100 were used to replace batteries in cars being used by the companys traveling sales staff. The remaining 6,300 batteries withdrawn from the storeroom were placed in cars being produced by the company. Of the cars in production during April, 90 percent were completec and transferred from work in process to finished goods. Of the cars completed during the month, 30 percent were unsold at April 30th Required . Determine the cost of batteries that would appear in each of the following accounts on April 30th. Name of the Account Cost 1a. Raw Materials 1b. Work in Process 1c. Finished Goods 1d. Cost of Goods Sold 1e. Selling Expense
Answer #1

Summary

Automobile Purchased

6,950

Cost Per unit

125

Automobile withdrawn from the storerooms

6400

Automobiles used in sales staff

100

Automobile used in production

6300

Completed and transferred to finished goods

90%

Unsold

30%

1.

Automobile Purchased

6,950

Less: withdrawn from the storerooms

6400

Automobile remaining in inventory

550

Cost Per Unit

125

Cost in raw mateials 550*125

68750

2. Work in Process

Automobiles used in production 6400 – 100

6300

Less: Completed and transferred to finished goods

90% * 6300 =

5670

Still in Work- in - Process

630

Cost per unit

125

Cost of work in process = 630 *125

78,750

3. Finished Goods

Completed and transferred to finished goods

90% * 6300 =

5670

Less: sold during the month 70%*6300

4410

Unsold in Finished goods

1260

Cost per unit

125

Cost of finished goods = 125 *1260

157,500

4. Cost of goods sold = Beginning inventory + Purchases – Ending Inventory

Purchases

6950

Less Ending inventory

1260

Cost of goods sold

5690

Cost of goods sold = 125 *5690

711,250

              

5. Selling Expense

sold during the month 70%*6300

4410

Cost per unit

125

4410 *125

551,250